# A classical density functional theory for solvation across length scales

### v2

@@ Abstract
A central aim of multiscale modeling is to use results from the Schrödinger Equation to predict phenomenology on length scales that far exceed those of typical molecular correlations. In this work, we present a new approach rooted in classical density functional theory (cDFT) that allows us to accurately describe the solvation of apolar solutes across length scales. Our approach builds on the Lum, Chandler and Weeks (LCW) theory of hydrophobicity [J. Phys. Chem. B 103, 4570 (1999)] by constructing a free energy functional that uses a slowly-varying component of the density field as a reference. From a practical viewpoint, the theory we present is numerically simpler and generalizes to solutes with soft-core repulsion more easily than LCW theory. Furthermore, by assessing the local compressibility and its critical scaling behavior, we demonstrate that our LCW-style cDFT approach contains the physics of critical drying, which has been emphasized as an essential aspect of hydrophobicity by recent theories. As our approach is parameterized on the two-body direct correlation function of the uniform fluid and the liquid–vapor surface tension, it straightforwardly captures the temperature dependence of solvation. Moreover, we use our theory to describe solvation at a first-principles level, on length scales that vastly exceed what is accessible to molecular simulations.
@@
$c_{\rm i}^{(1)}(\bm{\mathit{r}})=-\beta\frac{\delta\mathcal{F}^{(\rm ex)}_{\rm intr}}{\delta\rho(\bm{\mathit{r}})}\Bigg{|}_{{\rho_{\rm i}}},$ (3)
@@
Neglecting $\mathcal{F}_{\rm bridge}$ amounts to the hypernetted-chain approximation (HNCA) of integral equation theories [28]. Furthermore, upon linearizing $\Delta_{\rm u}\mathcal{F}_{\rm intr}^{\rm(id)}$, mDFT within the HNCA is equivalent, up to a small correction factor, to Chandler’s Gaussian field theory for solvation [29, 30]. Aside from being numerically tractable, the HNCA provides reasonable accuracy for small solutes. However, as we have emphasized, solvation is a multiscale problem. To see this, consider that the solute simply excludes solvent density within a radius $R$ of its center. Whereas within the HNCA, $\varOmega_{\rm solv}$ scales indefinitely with the solute’s volume, for large enough solutes we know from macroscopic theory that it scales with the solute’s surface area, i.e., $\varOmega_{\rm solv}\sim 4\pi\gamma R^{2}$ where $\gamma$ is the liquid–vapor surface tension. Many previous studies have shown that for water under ambient conditions, this “hydrophobic crossover” occurs at $R\approx 1$ nm [15, 31, 32, 33, 34].
The failure of the HNCA for large solutes arises because it cannot describe water’s proximity to its liquid–vapor coexistence at ambient conditions [35, 23]. A natural progression, then, is to attempt to encode the physics of coexistence through $\mathcal{F}_{\rm bridge}$, while maintaining reference to the homogeneous fluid with density $\rho_{\rm u}$. Such an approach has been used with some success within the mDFT framework for simple point charge water models [20, 21]. However, as we show in the S.I., existing bridge functionals of this kind are generally not robust to the choice of water model. In this article, we will instead follow more closely the key idea underlying LCW theory: the fluid’s density can be separated into a slowly-varying component that can sustain interfaces and liquid–vapor coexistence, and a rapidly-varying component that describes the local structure on microscopic length scales [15, 36].
In the following, we will describe how these ideas from LCW theory can be used to develop a cDFT approach to solvation of apolar solutes across both small and large length scales. We will validate our theory against available simulation data for the solvation of both hard- and soft-core solutes. We will show that the physics of critical drying, which is essential for a faithful description of solvophobicity on large length scales, is well-described. We will also demonstrate that our approach captures a distinguishing feature of solvophobicity in complex liquids such as water— the “entropic crossover”—that is absent in simple liquids. We will use our theory to describe the hydrophobic effect at an ab initio level, on length scales inaccessible to molecular simulations.
@@
In the HNCA the uniform fluid is assumed to act as a suitable reference density. In principle, we can perform a similar procedure where, instead of choosing a fluid of uniform density, we suppose that there exists some inhomogeneous, but slowly-varying, density field $\rho_{\rm s}(\bm{\mathit{r}})$ that acts as a suitable reference. In line with Eq. 1, $\rho_{\rm s}$ minimizes the grand potential functional $\varOmega_{{\phi_{\rm s}}}$ prescribed by a slowly-varying external potential $\phi_{\mathrm{s}}(\bm{\mathit{r}})$. Performing the expansion around $\rho_{\rm s}$ gives,
@@
Prescribing a general form for $\phi_{\rm s}$ is challenging. As we are interested in liquid water at thermodynamic states close to coexistence, we therefore adopt a strategy which exploits the fact that, at coexistence, inhomogeneous density fields minimize the grand potential when subject to appropriate boundary conditions. For example, in a planar geometry, solutions corresponding to the free liquid-vapor interface are obtained by minimizing the grand potential at coexistence, subject to the conditions $\rho(z\to-\infty)=\rho_{\rm v}$ and $\rho(z\to\infty)=\rho_{\rm l}$, where $\rho_{\rm v}$ and $\rho_{\rm l}$ are the vapor and liquid densities, respectively, at coexistence.
It is tempting to try and use such coexistence solutions directly as the reference density $\rho_{\rm s}$. However, the above example demonstrates the associated challenges. Imagine a liquid in contact with a planar hard wall. Close to coexistence, the average solvent density profile will resemble that of a free-liquid vapor interface with a well-defined separation from the wall. In contrast, in the above example, any density profile corresponding to free translation of the interface is equally plausible. We therefore seek a procedure that allows us to “pick” the appropriate coexistence solution. In fact, with the approximations that we will make, the coexistence solutions will ultimately only enter implicitly in our approach. To this end, we introduce a second expansion around an “auxiliary” reference density, $\rho_{\rm r}$, which is an equilibrium solution at coexistence. Expanding the slowly-varying reference density $\rho_{\rm s}$ around $\rho_{\rm r}$, we write
@@
In the context of solvation, how to choose boundary conditions on $\rho_{\rm r}$ is not immediately obvious. Moreover, we appear to have complicated matters, as we now need to deal with three density fields ($\rho$, $\rho_{\rm s}$ and $\rho_{\rm r}$). Assuming that the appropriate boundary conditions are specified, the first simplifying approximation that we make, which is reasonable near coexistence, is
@@
We also need a procedure for specifying $\rho_{\rm s}$ that keeps it sufficiently close to $\rho$ while maintaining a slowly-varying nature. To this end, we introduce the following “pseudofunctional,”
@@
The potential $\psi$ depends parametrically on both $\rho(\bm{\mathit{r}})$ and $\rho_{\rm s}(\bm{\mathit{r}})$, and penalizes differences between $\rho$ and $\rho_{\rm s}$, while ensuring that $\rho_{\rm s}$ remains slowly-varying. (We will specify a form for $\psi$ below.) In this case, $\mathcal{F}_{\mathrm{intr}}$ can be reasonably approximated by the square gradient form
@@
Equations 15, 18 and 19 provide the formal equations of our approach for the equilibrium density and grand potential. However, for practical calculations a form for the potential $\psi(\bm{\mathit{r}};[\rho,\rho_{\rm s}])$ needs to be specified. The fact that $\psi$ does not enter the grand potential explicitly in principle offers a great deal of flexibility, though as mentioned previously, it ought to penalize differences in $\rho(\bm{\mathit{r}})$ and $\rho(\bm{\mathit{r}})$ while ensuring that $\rho_{\rm s}(\bm{\mathit{r}})$ remains slowly-varying. With these considerations in mind, our choice for $\psi$ is guided by the final term in Eq. 12. Specifically, we assume that $c_{\rm r}^{(2)}$ can be separated as,
@@
$\psi(\bm{\mathit{r}})=-k_{\rm B}T\int\!\!\mathrm{d}\bm{\mathit{r}}^{\prime}\,c_{{\rm r},1}^{(2)}(\bm{\mathit{r}},\bm{\mathit{r}}^{\prime})[\rho(\bm{\mathit{r}})-\rho_{\rm s}(\bm{\mathit{r}}^{\prime})].$ (21)
@@
In Eq. 22, we see that the slowly-varying component of the two-body direct correlation function provides a natural means to coarse-grain the density fields. However, it is cumbersome to evaluate. Expressing $\psi$ in terms of the coarse-grained density fields as in Eq. 23 therefore also serves a practical purpose, as we can estimate its value with an approximate coarse-graining, e.g.,
$\overline{\rho}(\bm{\mathit{r}})\approx(2\pi\lambda^{2})^{-3/2}\int\!\!\mathrm{d}\bm{\mathit{r}}^{\prime}\,\exp(-|\bm{\mathit{r}}-\bm{\mathit{r}}^{\prime}|^{2}/2\lambda^{2})\rho(\bm{\mathit{r}}^{\prime}).$ (25)
@@
where $h(\rho_{\rm s})$ is the Heaviside step function. The curvature at the minima is determined by $C$, whose value we set to be consistent with the compressibility of the bulk fluid $\chi_{\rm u}$. Together, $D$ and $m$ determine $\gamma$ and the shape of the free liquid-vapor interface. Details of our parameterization procedure are provided in the S.I. In most of what follows we will treat $m$ as a constant independent of density; for spherical solutes, this yields the correct limiting behavior for both $R\to 0$ and $R\to\infty$. For $R$ in the crossover regime, we will show that the extra flexibility afforded by allowing $m$ to vary in a systematic, yet practical, way with $\rho_{\rm s}$ yields quantitative agreement for intermediate solute sizes.
@@
We first consider the paradigmatic test case for any theory of hydrophobicity: the solvation of a hard-sphere solute in water under ambient conditions ($T=300\,\mathrm{K}$, $\beta\delta\mu\approx 10^{-3}$). To this end, in Figs. 2(a) and 2(b) we present $\rho(r)$ and $\varOmega_{\rm solv}$ obtained from self-consistent solution of Eqs. 15 and 23, parameterized for a simple point charge model of water (SPC/E [39]), as we increase the solute radius $R$. First treating $m$ as a constant (dotted line in Fig. 2), we see that our cDFT approach broadly captures the behavior observed in molecular simulations [31], with $\varOmega_{\rm solv}$ increasing with volume for small solutes, before crossing over to $\varOmega_{\rm solv}/4\pi R^{2}\sim\gamma$ for $R\gtrsim 1$ nm. Like previous treatments rooted in LCW theory, however, we also see that these results overestimate $\varOmega_{\rm solv}$ in the crossover regime.
@@
Empirically, agreement with simulation results in the crossover regime could be attained by reducing the value of $m$. In the context of our theory, this implies that the optimal reference density for finite $R$ has a sharper interface than the free liquid-vapor interface. Simply reducing $m$ in such a manner, however, would result in an incorrect limiting behavior as $R\to\infty$. But this observation motivates us to vary $m$, in a linear fashion, with a characteristic feature of the slowly-varying density $\rho^{\ast}_{\rm s}$. Specifically, $\rho^{\ast}_{\rm s}$ is the value where $\rho_{\rm s}=\rho$ for the first time as $r$ increases. Further details are provided in the S.I. As we see in Fig. 2(b) (solid line), with $m(\rho^{\ast}_{\rm s})$ our cDFT describes the simulation data quantitatively.
@@
For hard spheres in water, the main advantage of our approach compared to LCW theory is primarily conceptual: $\varOmega_{\rm solv}$ follows directly from the minimization of $\varOmega_{\phi}$, as opposed to relying on, e.g., thermodynamic integration [40, 32, 41]. From a practical viewpoint, the cDFT approach is also numerically simpler to implement, as discussed in detail in Ref. 29. But its main advantages compared to LCW theory become apparent once we depart from the solvation of ideal hydrophobes. In LCW theory, it is assumed that the solute–solvent interaction can be separated into repulsive and attractive contributions, $\phi=\phi_{\mathrm{rep}}+\phi_{\mathrm{att}}$. While $\phi_{\mathrm{att}}$ can be accounted for straightforwardly in the slowly-varying part, $\phi_{\mathrm{rep}}$ must be approximated by a hard-core potential, and the solvent density is solved subject to the constraint $\rho(r)=0$ inside the solute [30]. For cDFT approaches such as ours, such an approximation is not needed. Instead, we simply minimize $\varOmega_{{\phi}}$ with the appropriate solute–solvent interaction $\phi$. To illustrate this, in Figs. 2(c) and 2(d), we show results from our theory for the solvation of Lennard–Jones solutes of effective radius $R$ (see S.I.) with a constant well-depth of $\epsilon=0.5\,\mathrm{kJ\,mol^{-1}}$. Using the same parameterization for $m(\rho^{\ast}_{\rm s})$ as for hard-spheres, our results for $\varOmega_{\rm solv}$ are in good agreement with available simulation data [38].
@@
which in practice is obtained from a finite-difference approximation (see S.I.). Evans, Wilding and co-workers argue that the structure of $\chi(\bm{\mathit{r}})$ provides one of the most robust indicators of hydrophobicity or, more generally, “solvophobicity.”
Figure 3: The local compressibility described by an LCW-style cDFT. In (a), we show $\chi(r)$ around different sized hard-sphere solutes, as indicated by the values of $R$ in the legend, centered at the origin. The results from our theory (left) are in good agreement with GCMC simulations from Ref. 47 (right). The inset shows $\chi_{\rm s}(r)$, the contribution from the slowly-varying density, which becomes increasingly important as $R$ increases, though contributions from the rapidly varying part are still significant. All results are for the mW water model ($T=426$ K, $\beta\delta\mu\approx 10^{-3}$, $a=300$ kJ cm3 mol-2 and $\lambda=0.11$ nm). In (b) we investigate the behavior of $\chi$ near a planar hard wall as $\delta\mu\to 0$. Consistent with a binding potential analysis [49], our theory yields $\ln\chi(\ell_{\rm eq})\sim-\ln\beta\delta\mu$ (right), where $\ell_{\rm eq}$ is the distance of the maximum in $\chi$ from the wall. We also observe $\ell_{\rm eq}\sim-\ln\beta\delta\mu$ (left). Circles show results from the theory and dashed lines indicate the expected scaling relation.
@@
A more stringent test of critical drying comes not from comparison to molecular simulations, but from known scaling behaviors of $\chi$ from binding potential analyses [49]. Specifically, for a fluid in contact with a planar hard wall, it can be shown that, close to coexistence, $\chi(\ell_{\rm eq})\sim\delta\mu^{-1}$, where $\ell_{\rm eq}$ is the position of the maximum in $\chi$. Moreover, $\ell_{\rm eq}\sim-\ln\delta\mu$. As seen in Fig. 3(b), the LCW-style cDFT approach that we have derived obeys both of these scaling relations. This is far from a trivial result. At face value, the physics of critical drying and the emphasis placed on liquid–vapor interface formation by LCW theory seem unrelated. By recasting the essential underpinnings of LCW theory in the context of cDFT, we begin to paint a unifying picture of these two different views on hydrophobicity.
@@
For illustrative purposes, we model water with RPBE-D3, a generalized gradient approximation (electronic) functional [54] with dispersion corrections [55]. To obtain $c^{(2)}_{\rm u}$, we have performed our own simulations of bulk water at the liquid density along the coexistence curve, using the MLIP described in Ref. 56, which also provides $\gamma(T)$. In Fig. 4(a), we present results from self-consistent solutions of Eqs. 15 and 23 for temperatures ranging from 300 K to 550 K. Note that, in the absence of information in the crossover regime, we have simply taken $m$ to be independent of density. For all temperatures, we observe the hydrophobic crossover. Importantly, this occurs at progressively smaller values of $R$ as $T$ increases; we observe the entropic crossover. These results demonstrate that the LCW-style cDFT that we have developed can be used to faithfully coarse-grain solvent effects while maintaining essential molecular correlations to model phenomenology at a first principles level across a broad range of length scales. This approach is summarized schematically in Fig. 4(b).
@@
The theory that we have outlined is similar in spirit to, and indeed motivated by, the seminal work on the hydrophobic effect by Lum, Chandler and Weeks [15]. By placing the ideas of LCW theory in the context of cDFT, not only do we gain a numerical advantage, but we also provide conceptual insights. For example, the grand potential that we derive (Eq. 12) suggests a natural form for the “unbalancing potential” that specifies the coarse-graining; the coarse-graining function is the slowly-varying part of a two-body direct correlation function of an inhomogeneous density field. This insight, as we explore in the S.I., justifies a coarse-graining length scale much smaller than the molecular diameter of a water molecule that one might naively expect.  [^2] For a Lennard–Jones fluid, this relationship between the coarse-grained density and the correlation function is identical to that specified by Weeks in the context of local molecular field theory [36], if we use a random phase approximation for the slowly-varying functional $\varOmega_{\rm s}$. However, it is clear that the overall functional that we obtain is not a simple mean field approximation.

### v3

@@ Abstract
A central aim of multiscale modeling is to use results from the Schrödinger equation to predict phenomenology on length scales that far exceed those of typical molecular correlations. In this work, we present a new approach rooted in classical density functional theory (cDFT) that allows us to accurately describe the solvation of apolar solutes across length scales. Our approach builds on the Lum, Chandler and Weeks (LCW) theory of hydrophobicity [K. Lum et al., J. Phys. Chem. B 103, 4570 (1999)] by constructing a free energy functional that uses a slowly-varying component of the density field as a reference. From a practical viewpoint, the theory we present is numerically simpler and generalizes to solutes with soft-core repulsion more easily than LCW theory. Furthermore, by assessing the local compressibility and its critical scaling behavior, we demonstrate that our LCW-style cDFT approach contains the physics of critical drying, which has been emphasized as an essential aspect of hydrophobicity by recent theories. As our approach is parameterized on the two-body direct correlation function of the uniform fluid and the liquid–vapor surface tension, it straightforwardly captures the temperature dependence of solvation. Moreover, we use our theory to describe solvation at a first-principles level, on length scales that vastly exceed what is accessible to molecular simulations.
Introduction
@@
$c_{i}^{(1)}(\bm{\mathit{r}})=-\beta\frac{\delta\mathcal{F}^{(\rm ex)}_{\rm intr}}{\delta\rho(\bm{\mathit{r}})}\Bigg{|}_{{\rho_{i}}},$ (3)
@@
Neglecting $\mathcal{F}_{\rm bridge}$ amounts to the hypernetted-chain approximation (HNCA) of integral equation theories [28]. Furthermore, upon linearizing $\Delta_{\rm u}\mathcal{F}_{\rm intr}^{\rm(id)}$, mDFT within the HNCA is equivalent, up to a small correction factor, to Chandler’s Gaussian field theory for solvation [29, 30]. Aside from being numerically tractable, the HNCA provides reasonable accuracy for small solutes. However, as we have emphasized, solvation is a multiscale problem. To see this, consider that the solute simply excludes solvent density within a radius $R$ of its center. Whereas within the HNCA, $\varOmega_{\rm solv}$ scales indefinitely with the solute’s volume [30, 31, 29, 32], for large enough solutes we know from macroscopic theory that it scales with the solute’s surface area, i.e., $\varOmega_{\rm solv}\sim 4\pi\gamma R^{2}$ where $\gamma$ is the liquid–vapor surface tension. Many previous studies have shown that for water under ambient conditions, this “hydrophobic crossover” occurs at $R\approx 1$ nm [15, 33, 34, 35, 36].
The failure of the HNCA for large solutes arises because it cannot describe water’s proximity to its liquid–vapor coexistence at ambient conditions [37, 23]. A natural progression, then, is to attempt to encode the physics of coexistence through $\mathcal{F}_{\rm bridge}$, while maintaining reference to the homogeneous fluid with density $\rho_{\rm u}$. Such an approach has been used with some success within the mDFT framework for simple point charge water models [20, 21]. However, as we show in the Supplementary Material (SM), existing bridge functionals of this kind are generally not robust to the choice of water model. We also note that attempts to use the “hard-sphere bridge functional” for water [38, 39, 18] have proved problematic, failing to describe simultaneously the pressure and surface tension [20]. In this article, we will instead follow more closely the key idea underlying LCW theory: the fluid’s density can be separated into a slowly-varying component that can sustain interfaces and liquid–vapor coexistence, and a rapidly-varying component that describes the local structure on microscopic length scales [15, 40].
In the following, we will outline how these ideas from LCW theory can be used to develop a cDFT approach to describe the solvation of apolar solutes across both small and large length scales. We will validate our theory against available simulation data for the solvation of both hard- and soft-core solutes. We will show that the physics of critical drying, which is essential for a faithful description of solvophobicity on large length scales, is well-described. We will also demonstrate that our approach captures a distinguishing feature of solvophobicity in complex liquids such as water— the “entropic crossover”—that is absent in simple liquids. We will use our theory to describe the hydrophobic effect at an ab initio level, on length scales inaccessible to molecular simulations.
@@
In the HNCA, the uniform fluid is assumed to act as a suitable reference density. In principle, we can perform a similar procedure where, instead of choosing a fluid of uniform density, we suppose that there exists some inhomogeneous, but slowly-varying, density field $\rho_{\rm s}(\bm{\mathit{r}})$ that acts as a suitable reference. In line with Eq. 1, $\rho_{\rm s}$ minimizes the grand potential functional $\varOmega_{{\phi_{\rm s}}}$ prescribed by a slowly-varying external potential $\phi_{\mathrm{s}}(\bm{\mathit{r}})$. Performing the expansion around $\rho_{\rm s}$ gives,
@@
Prescribing a general form for $\phi_{\rm s}$ is challenging. As we are interested in liquid water at thermodynamic states close to coexistence, we therefore adopt a strategy which exploits the fact that, at coexistence, inhomogeneous density fields minimize the grand potential when subject to appropriate boundary conditions. For example, in a planar geometry, solutions corresponding to the free liquid-vapor interface are obtained by minimizing the grand potential at coexistence, subject to the conditions
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho(z\to-\infty)=\rho_{\rm v}\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }\rho(z\to\infty)=\rho_{\rm l},}$
where $\rho_{\rm v}$ and $\rho_{\rm l}$ are the vapor and liquid densities, respectively, at coexistence.
It is tempting to try and use such coexistence solutions directly as the reference density $\rho_{\rm s}$. However, the above example demonstrates the associated challenges. Imagine a liquid in contact with a planar hard wall. Close to coexistence, the average solvent density profile will resemble that of a free liquid–vapor interface with a well-defined separation from the wall. In contrast, in the above example, any density profile corresponding to free translation of the interface is equally plausible. We therefore seek a procedure that allows us to “pick” the appropriate coexistence solution. In fact, with the approximations that we will make, the coexistence solutions will ultimately only enter implicitly in our approach. To this end, we introduce a second expansion around an “auxiliary” reference density, $\rho_{\rm r}$, which is an equilibrium solution at coexistence. Expanding the slowly-varying reference density $\rho_{\rm s}$ around $\rho_{\rm r}$, we write
@@
In the context of solvation, how to choose boundary conditions on $\rho_{\rm r}$ is not immediately obvious. Moreover, we appear to have complicated matters, as we now need to deal with three density fields ($\rho$, $\rho_{\rm s}$ and $\rho_{\rm r}$). We will now outline a series of approximations that enable us to consider a theory expressed explicitly in terms of $\rho$ and $\rho_{\rm s}$ only, as well as a procedure to “pick” an appropriate slowly-varying reference density.
The first simplifying approximation that we make, which is reasonable near coexistence, is
@@
To construct a procedure for specifying $\rho_{\rm s}$ that keeps it sufficiently close to $\rho$ while maintaining a slowly-varying nature, we introduce the following “pseudofunctional,”
@@
We use the term “pseudofunctional” because $\psi$ is not an external potential; it depends parametrically on both $\rho(\bm{\mathit{r}})$ and $\rho_{\rm s}(\bm{\mathit{r}})$, and penalizes differences between $\rho$ and $\rho_{\rm s}$, while ensuring that $\rho_{\rm s}$ remains slowly-varying. (We will specify a form for $\psi$ below.) In this sense, it is not a functional in the usual cDFT sense, and we have introduced it simply as a computational tool to obtain an appropriate slowly-varying reference.
As $\rho_{\rm s}$ is slowly-varying, $\mathcal{F}_{\mathrm{intr}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[\rho_{\rm s}]}$ can be reasonably approximated by the square gradient form
@@
Equations 15, 18 and 19 provide the formal equations of our approach for the equilibrium density and grand potential. However, for practical calculations, a form for the potential $\psi(\bm{\mathit{r}};[\rho,\rho_{\rm s}])$ needs to be specified. In this work, we will not explore ways to systematically optimize $\psi$. While, in principle, the fact that $\psi$ does not enter the grand potential explicitly offers some flexibility in specifying its form, it is not completely arbitrary; as mentioned previously, it ought to penalize differences between $\rho(\bm{\mathit{r}})$ and $\rho_{\rm s}(\bm{\mathit{r}})$ while ensuring that $\rho_{\rm s}(\bm{\mathit{r}})$ remains slowly-varying. (Also see below, where we discuss our approach in the context of LCW theory.) With these considerations in mind, our choice for $\psi$ is guided by the final term in Eq. 12. Specifically, we assume that $c_{\rm r}^{(2)}$ can be separated as,
@@
$\psi(\bm{\mathit{r}})=-k_{\rm B}T\int\!\!\mathrm{d}\bm{\mathit{r}}^{\prime}\,c_{{\rm r},1}^{(2)}(\bm{\mathit{r}},\bm{\mathit{r}}^{\prime})[\rho(\bm{\mathit{r}}^{\prime})-\rho_{\rm s}(\bm{\mathit{r}}^{\prime})].$ (21)
@@
Note that, if we were interested in a Lennard–Jones fluid, this relationship between the coarse-grained density and the correlation function is identical to that specified by Weeks in the context of local molecular field theory [40], in which $c^{(2)}_{{\rm r},1}$ is treated in a random phase approximation. (Although also note that Eqs. 15, 18 and 19 do not constitute a random phase approximation.) In this case, there is a notion of “unbalanced attractive interactions,” and in the original LCW theory [15], it was argued that such unbalanced forces in water could be described by a similar coarse graining, over an appropriate length scale that describes the range of attractive interactions.
Our cDFT-based approach does not emphasise the role of unbalanced attractive forces. In Eq. 22, we see that the slowly-varying component of the two-body direct correlation function provides a natural means to coarse-grain the density fields. However, it is cumbersome to evaluate. Expressing $\psi$ in terms of the coarse-grained density fields as in Eq. 23 therefore also serves a practical purpose, as we can estimate its value with an approximate coarse-graining, e.g.,
$\overline{\rho}(\bm{\mathit{r}})\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{(2\pi\lambda^{2})^{3/2}}}\int\!\!\mathrm{d}\bm{\mathit{r}}^{\prime}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp\left(-\frac{|\bm{\mathit{r}}-\bm{\mathit{r}}^{\prime}|^{2}}{2\lambda^{2}}\right)}\rho(\bm{\mathit{r}}^{\prime}).$ (25)
@@
where $h(\rho_{\rm s})$ is the Heaviside step function. The curvature at the minima is determined by $C$, whose value we set to be consistent with the compressibility of the bulk fluid $\chi_{\rm u}$. Together, $D$ and $m$ determine $\gamma$ and the shape of the free liquid–vapor interface. Details of our parameterization procedure are provided in the SM. In most of what follows we will treat $m$ as a constant independent of density; for spherical solutes, this yields the correct limiting behavior for both $R\to 0$ and $R\to\infty$. For $R$ in the crossover regime, we will show that the extra flexibility afforded by allowing $m$ to vary in a systematic, yet practical, way with $\rho_{\rm s}$ yields quantitative agreement for intermediate solute sizes.
@@
We first consider the paradigmatic test case for any theory of hydrophobicity: the solvation of a hard-sphere solute in water under ambient conditions ($T=300\,\mathrm{K}$, $\beta\delta\mu\approx 10^{-3}$). To this end, in Figs. 2(a) and 2(b) we present $\rho(r)$ and $\varOmega_{\rm solv}$ obtained from self-consistent solutions of Eqs. 15 and 23, parameterized for a simple point charge model of water (SPC/E [43]), as we increase the solute radius $R$. First treating $m$ as a constant (dotted line in Fig. 2), we see that our cDFT approach broadly captures the behavior observed in molecular simulations [33], with $\varOmega_{\rm solv}$ increasing with volume for small solutes, before crossing over to $\varOmega_{\rm solv}/4\pi R^{2}\sim\gamma$ for $R\gtrsim 1$ nm. Like previous treatments rooted in LCW theory, however, we also see that these results overestimate $\varOmega_{\rm solv}$ in the crossover regime.
@@
Empirically, agreement with simulation results in the crossover regime could be attained by reducing the value of $m$. In the context of our theory, this implies that the optimal reference density for finite $R$ has a sharper interface than the free liquid-vapor interface. Simply reducing $m$ in such a manner, however, would result in an incorrect limiting behavior as $R\to\infty$. But this observation motivates us to vary $m$, in a linear fashion, with a characteristic feature of the slowly-varying density $\rho^{\ast}_{\rm s}$. Specifically, $\rho^{\ast}_{\rm s}$ is the value where $\rho_{\rm s}=\rho$ for the first time as $r$ increases. Further details are provided in the SM. As we see in Fig. 2(b) (solid line), with $m(\rho^{\ast}_{\rm s})$ our cDFT describes the simulation data quantitatively.
@@
For hard spheres in water, the main advantage of our approach compared to LCW theory is primarily conceptual: $\varOmega_{\rm solv}$ follows directly from the minimization of $\varOmega_{\phi}$, as opposed to relying on, e.g., thermodynamic integration [44, 34, 45]. From a practical viewpoint, the cDFT approach is also numerically simpler to implement, as discussed in detail in Ref. 29. But its main advantages compared to LCW theory become apparent once we depart from the solvation of ideal hydrophobes. In LCW theory, it is assumed that the solute–solvent interaction can be separated into repulsive and attractive contributions, $\phi=\phi_{\mathrm{rep}}+\phi_{\mathrm{att}}$. While $\phi_{\mathrm{att}}$ can be accounted for straightforwardly in the slowly-varying part, $\phi_{\mathrm{rep}}$ must be approximated by a hard-core potential, and the solvent density is solved subject to the constraint $\rho(r)=0$ inside the solute [30]. For cDFT approaches such as ours, such an approximation is not needed. Instead, we simply minimize $\varOmega_{{\phi}}$ with the appropriate solute–solvent interaction $\phi$. To illustrate this, in Figs. 2(c) and 2(d), we show results from our theory for the solvation of Lennard–Jones solutes of effective radius $R$ (see SM) with a constant well-depth of $\epsilon=0.5\,\mathrm{kJ\,mol^{-1}}$. Using the same parameterization for $m(\rho^{\ast}_{\rm s})$ as for hard-spheres, our results for $\varOmega_{\rm solv}$ are in good agreement with available simulation data [42].
@@
which in practice is obtained from a finite-difference approximation (see SM). Evans, Wilding and co-workers argue that the structure of $\chi(\bm{\mathit{r}})$ provides one of the most robust indicators of hydrophobicity or, more generally, “solvophobicity.”
Figure 3: The local compressibility described by an LCW-style cDFT. In (a), we show $\chi(r)$ around different sized hard-sphere solutes, as indicated by the values of $R$ in the legend, centered at the origin. The results from our theory (left) are in good agreement with GCMC simulations from Ref. 51 (right). The inset shows $\chi_{\rm s}(r)$, the contribution from the slowly-varying density, which becomes increasingly important as $R$ increases, though contributions from the rapidly varying part are still significant. All results are for the mW water model ($T=426$ K, $\beta\delta\mu\approx 10^{-3}$, $a=300$ kJ cm3 mol-2 and $\lambda=0.11$ nm). In (b) we investigate the behavior of $\chi$ near a planar hard wall as $\delta\mu\to 0$. Consistent with a binding potential analysis [53], our theory yields $\ln\chi(\ell_{\rm eq})\sim-\ln\beta\delta\mu$ (right), where $\ell_{\rm eq}$ is the distance of the maximum in $\chi$ from the wall. We also observe $\ell_{\rm eq}\sim-\ln\beta\delta\mu$ (left). Circles show results from the theory and dashed lines indicate the expected scaling relation.
@@
A more stringent test of critical drying comes not from comparison to molecular simulations, but from known scaling behaviors of $\chi$ from binding potential analyses [53]. Specifically, for a fluid in contact with a planar hard wall, it can be shown that, close to coexistence, $\chi(\ell_{\rm eq})\sim\delta\mu^{-1}$, where $\ell_{\rm eq}$ is the position of the maximum in $\chi$. Moreover, $\ell_{\rm eq}\sim-\ln{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\delta\mu$. As seen in Fig. 3(b), the LCW-style cDFT approach that we have derived obeys both of these scaling relations. This is far from a trivial result. At face value, the physics of critical drying and the emphasis placed on liquid–vapor interface formation by LCW theory seem unrelated. By recasting the essential underpinnings of LCW theory in the context of cDFT, we begin to paint a unifying picture of these two different views on hydrophobicity.
@@
For illustrative purposes, we model water with RPBE-D3, a generalized gradient approximation (electronic) functional [58] with dispersion corrections [59]. To obtain $c^{(2)}_{\rm u}$, we have performed our own simulations of bulk water at the liquid density along the coexistence curve, using the MLIP described in Ref. 60, which also provides $\gamma(T)$. In Fig. 4(a), we present results from self-consistent solutions of Eqs. 15 and 23 for temperatures ranging from 300 K to 550 K. Note that, in the absence of information in the crossover regime, we have simply taken $m$ to be independent of density. For all temperatures, we observe the hydrophobic crossover. Importantly, this occurs at progressively smaller values of $R$ as $T$ increases; we observe the entropic crossover. These results demonstrate that the LCW-style cDFT that we have developed can be used to faithfully coarse-grain solvent effects while maintaining essential molecular correlations to model phenomenology at a first principles level across a broad range of length scales. This approach is summarized schematically in Fig. 4(b).
@@
The theory that we have outlined is similar in spirit to, and indeed motivated by, the seminal work on the hydrophobic effect by Lum, Chandler and Weeks [15]. By placing the ideas of LCW theory in the context of cDFT, not only do we gain a numerical advantage, but we also provide conceptual insights. For example, the grand potential that we derive (Eq. 12) suggests a natural form for the “unbalancing potential” that specifies the coarse-graining; the coarse-graining function is the slowly-varying part of a two-body direct correlation function of an inhomogeneous density field. This insight, as we explore in the SM, justifies a coarse-graining length scale much smaller than the molecular diameter of a water molecule that one might naively expect.